OpenStudy (anonymous):

A solid wooden cylinder of height 8cm and radius 3cm is cut in two vertical axis of symmetry. Calculate the total surface area of the two pieces

OpenStudy (anonymous):

curved surface area = pi *3^2*8=72 pi area of the planar circular face = 3*3 pi (each) hence required total surface area =(72 pi)/4 + ((9 pi)/4 )*2 +2*8*3